OpenStudy (moonlitfate):

A population of bacteria is changing at the rate of dP/dt = 2000/(1+0.2t), where t is the time in days. The initial population when t=0 is 1000. Determine the population after 10 days.

OpenStudy (anonymous):

1000*log(1+0.2t), then put 10 into the new function.

OpenStudy (anonymous):

I get the answer 1439 dP=2000/(1+0.2t) dt integrate the right hand term from 0 to 10 left hand term, integrate from 1000 to P you can get the answer

OpenStudy (anonymous):

Note that we can integrate this to receive a non-linear expression. A hint: \[ \int\frac{dP}{dt}dt=2000\int\frac{1}{1+.2t}dt \]Which turns to: \[ \int dP=2000\int\frac{1}{1+.2t}dt \]Using u-substitution for \(t\), you'll receive the function, which you can use to evaluate what you're given.

OpenStudy (moonlitfate):

@LolWolf -- I never really understood the whole u-substitution thing; it just makes things more confusing, lol. Though it probably is supposed to do the opposite.

OpenStudy (anonymous):

use Woframalpha .com. your question is a integral question.

OpenStudy (anonymous):

All right, I'll explain what u-sub is, it's actually pretty intuitive, and kind of nice. Note the chain rule in differentiation: \[ \frac{d}{dx}f(g(x))=f'(g(x))g'(x) \]Now, integrate both sides of the expression, where we receive: \[ f(g(x))=\int f'(g(x))g'(x)\;dx \]But we also know, by definition: \[ f(g(x))=\int f'(g(x))\;dg \]So \[ f(g(x))=\int f'(g(x))\frac{dg}{dx}\;dx=\int f'(g(x))\;dg \]So this actually turns into: \[ \int f'(g(x))\frac{dg}{dx}\;dx=\int f'(g(x))\;dg=f(g(x)) \]Right? So, u-substitution is doing exactly that, we're changing the variable of integration based on another function.

OpenStudy (moonlitfate):

Oh, all right. So, let me see if I follow this:\[\int\limits_{}^{}\frac{ 1 }{ 1+.2t }dt\]f(t) = \[1*(1+.2t)^-1\] So, does that mean u would be the demonimator of the fraction? Sorry, bear with me if I'm wrong, trying to u nderstand.

OpenStudy (anonymous):

Yep, that works. So, we now have: \[ u=1+.2t \]So then: \[ du=\;? \]

OpenStudy (moonlitfate):

du = \[0+0.2 = 0.2\]

OpenStudy (anonymous):

Don't forget your dt. du=0.2dt But yes. you are correct.

OpenStudy (anonymous):

So, back to your original integral. You need to solve for dt. So, dt=du/0.2 =5du. So, your integral changes to...

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ 5du }{ u}\]

OpenStudy (moonlitfate):

Oh, right since dt was solved for; this is still making little sense, but I'm trying to understand, so. Would you have to plug the values back in for du and u?

OpenStudy (anonymous):

Well, the nice thing about using u-sub, is that this is now a very easy function to integrate. It integrates into 5* ln |u|. Now you can go back and replace what u is equal to.

OpenStudy (anonymous):

OF course don't forget your constant of integration so that you can evaluate your initial condition.

OpenStudy (moonlitfate):

so \[\int\limits_{}^{}\frac{ 5(0.2dt) }{ 5*\ln \left| 1+0.2t \right|}\]

OpenStudy (moonlitfate):

Oh, I wouldn't need dt in the numerator. Would just be 5(0.2), right?

OpenStudy (anonymous):

Well, not quite, the integral(anti-derivative) has been found, so we would remove the integral sign, as well as dropping the entire numerator. Your antiderivat of the entire u-sub is going to be 5* ln |1+0.2t| + C.

OpenStudy (anonymous):

Now, you have a 2000 sitting out in front of that integral, so you will need to multiply the result by 2000, giving your 10000*ln |1+0.2t| + C. Now, you must find C when t=0, then you will have your function for any time t.

OpenStudy (anonymous):

So, P(t)= 10000*ln|1+0.2t| + C, and you know P(0)=1000, so plug that information into your function and you should be able to determine C.

OpenStudy (anonymous):

Questions, confusion?

OpenStudy (moonlitfate):

Sorry, still working it out over here.

OpenStudy (anonymous):

ok.

OpenStudy (moonlitfate):

I got -9000.

OpenStudy (moonlitfate):

Oh, wait, wrong... lol.

OpenStudy (moonlitfate):

C=1000

OpenStudy (anonymous):

good

OpenStudy (anonymous):

Now, combine that with the first part and you have ur function for any time, t. So, all that is left is to find P(10).

OpenStudy (moonlitfate):

So, P(t) = \[10000*\ln \left| 1+0.2t \right|+1000\] P(10) = 10000*ln|1+0.2(10)|+1000

OpenStudy (anonymous):

perfect! what is that value?

OpenStudy (moonlitfate):

So P = 110861? :D

OpenStudy (anonymous):

hmmm...one too many decimals in that answer. Try again.

OpenStudy (moonlitfate):

11986.1

OpenStudy (anonymous):

There we go!

OpenStudy (moonlitfate):

Wow! Thank you so much. I appreciate all this help with math problems that I have little idea how to do. I would give you 1 million medals if I could.

OpenStudy (anonymous):

lol...nah, it's my job. For some reason I like to do it at work and at home.

OpenStudy (anonymous):

Good luck with the rest of your work. I will be on for a bit longer if you get stuck on another one.